# BSTcontrol

### v2

@@ 2.1.5 Diffusion Related Architectures
Denoising diffusion probabilistic models (DDPMs) are a class of generative models that create data samples through an iterative denoising process [46]. They work by gradually corrupting real data with noise during training, and then learning to reverse this process. The model is typically trained to predict either the original data or the noise added at various levels of degradation. The loss function is denoted as
$\mathcal{L}_{\text{DDPM}}=\mathbb{E}_{t,\mathbf{x}_{0},\bm{\epsilon}}\left[||\bm{\epsilon}-\bm{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t)||^{2}\right],$ (1)
@@
Denoising diffusion implicit models (DDIMs) are a variant of DDPMs that focus on improving generation efficiency [47]. While DDPMs rely on stochastic sampling, DDIMs use a deterministic reverse process to generate data, denoted as
@@
Diffusion transformers (DiTs) are used by transforming input images into a sequence of patches, which are then processed by transformer blocks [48]. To use DiTs, the images need to undergo a patchifying step, which breaks them into smaller tokens. These tokens are then passed through multiple DiT blocks, which use self-attention mechanisms and optional conditioning techniques (such as adaptive layer normalization) to handle various input factors like noise or class labels. DiTs can be trained with latent diffusion models to reduce computational cost, and they scale efficiently with increased model size, improving image quality as the model size or computational resources grow. DiT models like Hunyuan-DiT [49] have gained attention for their outstanding capability in processing pictures.
2.1.6 Mixture-of-Experts
Mixture-of-experts (MoE) is a scalable model architecture that dynamically routes inputs to different specialized subnetworks, known as experts [50]. A gating mechanism selects a subset of these experts based on the input, allowing only the chosen experts to process the data. This conditional computation enables models with massive parameter counts to operate efficiently, as only a small portion of the network is active at a time. MoE architectures have demonstrated impressive scalability, particularly in LLMs like DeepSeek-V3 [51], where they improve performance while maintaining inference cost. However, training such models with numerous experts can be challenging, particularly in ensuring the gating mechanism effectively learns to route inputs. In dynamic wireless environments, these challenges manifest as channel conditions and data distributions fluctuate. Recent advancements like DeepEP [52], DeepGEMM [53] addressed these issue by optimizing MoE training with efficient graphic processing unit (GPU) communication and accelerated matrix operations. In wireless applications, MoEs could enable adaptive processing by activating experts tailored to specific environments, such as urban or rural channel conditions, thereby enhancing model specialization and generalization.
@@ 2.2.1 Data Sources and Diversity
Pretraining large models requires diverse datasets. In wireless, this includes channel measurements, CSI matrices, beam logs, and traffic traces collected from real-world networks or high-fidelity simulators. Ensuring coverage across different frequency bands, mobility profiles, and deployment scenarios improves generalization. Unlike conventional pre-training that relies heavily on existing static datasets sourced from the internet, wireless pre-training leverages real-time, multi-modal data collected from a diverse array of internet of things (IoT) devices, including smartphones, sensors, and autonomous as well. This dynamic and heterogeneous data reflects a more accurate and timely representation of the physical world, enhancing the ability of models to understand and predict complex patterns and interactions in various environments. Synthetic data which is generated using ray-tracing, digital twins, or data augmentation strategies to simulate rare or extreme events have also gain consideration [54]. In semantic communication systems, data must also cover various modalities (text, image, video, audio) and tasks (translation, reconstruction, classification) to ensure the model can learn semantic compression and transmission strategies that generalize across use cases.
@@ 2.2.3 Ethical Considerations
Large-scale wireless datasets may inadvertently capture sensitive user patterns. Datasets often reflect existing social, geographic, or behavioral biases. Just as general-purpose language models may struggle with underrepresented languages or dialects, WLAM can exhibit performance gaps if trained predominantly on data from urban areas or specific hardware configurations. For example, overrepresenting urban environments or specific hardware configurations may result in models that underperform in rural or heterogeneous conditions. Ensuring balanced representation in training data across demographics, devices, environments, and behaviors is essential to avoid digital exclusion or uneven service quality. Besides, communication data is sensitive by its nature. IQ samples, packet logs, or location traces can inadvertently contain personal identifiers, usage patterns, or behavioral signals. Wireless systems are particularly vulnerable to misuse, as radio signals inherently propagate into shared environments. This makes it essential to align data collection with regulatory standards such as general data protection regulation [55] or California consumer privacy act [56].
@@ 2.3.1 Training Objectives
The choice of training objective plays a critical role in shaping what a model learns during pre-training, with different objectives aligning with specific model architectures and downstream use cases. The next-token prediction objective, commonly used in GPT-style decoder-only models, trains the model to predict the next element in a sequence given the preceding tokens [42]. This approach excels in generative tasks and is well-suited to sequence modeling problems in wireless systems, such as predicting channel evolution or traffic patterns. Masked modeling, used in encoder-only architectures like BERT, involves masking random portions of the input and training the model to reconstruct them using bidirectional context. In wireless applications, this method can be employed for tasks such as recovering missing signal samples or corrupted subcarriers. Sequence-to-sequence generation objectives, typical of encoder-decoder models like T5 or BART, transform an input sequence into a target output, which is particularly useful for end-to-end communication systems, semantic compression, or signal-to-action mappings in adaptive wireless control.
@@ 2.3.2 Distributed and Large-Scale Training
Large model training involves billions of parameters and requires substantial computational resources. Supervised pre-training is typically carried out over distributed systems that involve multiple graphics processing units (GPUs) connected through high-speed interconnects. To efficiently scale across devices, training workflows employ parallelization strategies such as data parallelism, where each device trains on a separate batch of data with synchronized gradients, or model parallelism, where the model itself is partitioned across devices to handle larger parameter sizes than a single GPU can accommodate.
In pipeline parallelism, the layers of the model are distributed across devices and trained in a staged fashion, enabling higher throughput. WLAM, particularly those involving long input sequences or multimodal inputs, benefit from such techniques. Mixed-precision training reduces memory usage and increases compute efficiency by using lower-precision arithmetic without sacrificing accuracy, and sharded state parallel [57] reduces the overhead of large optimizer memory footprints. Specially designed methods like Dualpipe [58] also provide an innovative bidirectional pipeline parallelism algorithm and implementation.
Figure 4:   Reinforcement Learning Techniques.
@@ 2.4 Reinforcement Learning
Reinforcement learning (RL) is a method where an AI agent learns by interacting with an environment, receiving rewards or penalties for its actions, and aiming to maximize long-term rewards. In the context of large AI models, RL is particularly valuable for tasks where direct supervision is hard to define, such as reasoning, alignment with human values, or optimizing outcomes over multiple steps. RL models these tasks as a markov decision process and uses a policy to select actions based on the current state, with the goal of maximizing expected cumulative rewards. RL shares similarities with adaptive wireless systems, where decisions like scheduling are made under uncertainty and delayed feedback. Core components include the policy, action space, reward function, and optionally a value estimator. Unlike supervised learning, which relies on static labels, RL enables dynamic exploration, making it well-suited for environments with non-stationary or ambiguous objectives, common in both wireless networks and large-scale model alignment, as summarized in Fig. 4. As large models scale, hybrid training strategies could further enhance efficiency. Moreover, by integrating RL with the existing supervised pre-training, models can adjust to the evolving demands of wireless environments.
2.4.1 Proximal Policy Optimization
Proximal policy optimization (PPO) [59] is a policy gradient RL algorithm that has gained significant popularity due to its stability and efficiency, making it a common choice for fine-tuning LLMs, especially in the context of RL from human feedback. PPO operates by collecting a set of trajectories through interaction with the environment using the current policy. It then computes the rewards obtained in these trajectories and estimates the advantage of each action taken. The policy is updated by maximizing a surrogate objective function that aims to improve the performance of the policy while ensuring that the updates to the policy are not excessively large. This constraint on the policy update is typically enforced using a clipping mechanism that limits the ratio between the probability of an action under the new policy and its probability under the old policy, denoted as
@@ 2.4.1 Proximal Policy Optimization
which represents the core PPO clipped surrogate objective function $L^{\text{CLIP}}(\theta)$. The expectation $\hat{\mathbb{E}}_{t}$ is over a batch of transitions. $r_{t}(\theta)=\pi_{\theta}(a_{t}|s_{t})/\pi_{\theta_{\text{old}}}(a_{t}|s_{t})$ is the probability ratio of action $a_{t}$ under the current policy $\pi_{\theta}$ versus the policy $\pi_{\theta_{\text{old}}}$ used for data collection. $\hat{A}_{t}$ is the estimated advantage of taking action $a_{t}$ in state $s_{t}$. The clipping mechanism (using hyperparameter $\epsilon$) restricts the policy change per update, enhancing stability. PPO seeks to strike a balance between exploration, which involves trying out new actions to discover potentially better strategies, and exploitation, which involves taking actions that are known to yield high rewards based on the current knowledge. PPO is widely used due to its simplicity, efficiency, and robustness when scaling to large models.
2.4.2 Reinforcement Learning From Human Feedback
Reinforcement learning from human feedback (RLHF) aligns LLMs with human preferences [60, 61]. It is especially effective in tasks where human judgment is clear but difficult to formalize. By incorporating human feedback, RLHF helps reduce biases and unsafe behavior, making it central to aligning models like InstructGPT [62] and OpenAI o1 with human values. Unlike supervised learning, RLHF incorporates human judgment, refining model behavior in tasks where objectives are hard to define. The RLHF pipeline typically involves three stages: supervised fine-tuning on instruction-following examples, training a reward model with human-labeled preferences, and optimizing the model with RL, often using PPO to ensure stability and efficiency, denoted as
@@ 2.4.2 Reinforcement Learning From Human Feedback
which shows the objective maximized during the RL phase of RLHF. The language model policy $\pi_{\theta}$ is optimized to generate responses $y$ for prompts $x$ (from distribution $\mathcal{D}$) that maximize the score from a learned reward model $R_{\phi}(x,y)$, which reflects human preferences. A KL-divergence penalty, weighted by $\beta$, regularizes the policy $\pi_{\theta}$ to stay close to a reference policy $\pi_{\text{ref}}$ (e.g., the initial SFT model), maintaining model capabilities and stability. On the other hand, RLHF faces challenges, including the need for high-quality human feedback, the complexity of reward model training, and computational intensity. Additionally, there are risks of reward hacking [63], where models exploit reward signals without genuinely improving outputs. Despite these challenges, RLHF remains a powerful method for scalable LLM alignment and has potential applications in WLAM.
@@
2.4.3 Direct Preference Optimization
Direct preference optimization (DPO) [64] is a alternative to RLHF that removes the need for an explicit reward model. DPO directly optimizes the language model using human preference data by framing alignment as a classification task. Given pairs of model outputs labeled by human preference, DPO adjusts model parameters to increase the likelihood of preferred responses, typically using a binary cross-entropy loss, as
@@ 2.4.3 Direct Preference Optimization
2.4.4 Group Relative Policy Optimization
@@
which defines the group-relative advantage estimate central to GRPO. For a prompt $x$, the policy generates a group of $K$ candidate responses $\{y_{j}\}_{j=1}^{K}$. The advantage $\hat{A}_{G}$ of a specific response $y_{i}$ is computed as its reward $R(x,y_{i})$ (obtained from a reward model or human feedback) minus the average reward across all $K$ responses in that group. This relative advantage measure replaces the value function estimate used in traditional policy gradient methods like PPO. This approach simplifies training, reduces memory usage, and eliminates the need for large-scale critic gradient updates. Specifically, in DeepSeek-R1-Zero [65], GRPO was applied without any supervised fine-tuning, yielding a model trained purely via RL. While early iterations struggled with readability, later versions like DeepSeek-R1 combined GRPO with multi-stage pretraining and alignment to achieve state-of-the-art results on reasoning tasks. The scalability of GRPO has also influenced other models such as QwQ-32B [66], which uses a GRPO-inspired two-stage RL method to achieve strong performance with just 32B parameters.

### v4

@@ 2.1.5 Diffusion Related Architectures
Basic Principle of Diffusion Models
Diffusion models have emerged as a leading paradigm in generative modeling, distinguished by their capacity for high-fidelity sample generation and stable training dynamics. The core principle is inspired by non-equilibrium thermodynamics, which describes how systems evolve from order to disorder. This concept is reflected in two opposing processes: a fixed forward diffusion process and a learned reverse denoising process. In the forward process, real data $\mathbf{x}_{0}$ is gradually perturbed by Gaussian noise through $T$ discrete steps:
$q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\sqrt{\alpha_{t}}\mathbf{x}_{t-1},(1-\alpha_{t})\mathbf{I}),$ (1)
where $\alpha_{t}\in(0,1)$ controls the noise level at step $t$, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ denotes the cumulative product.
By iterating this process, we obtain $\mathbf{x}_{T}\approx\mathcal{N}(\mathbf{0},\mathbf{I})$, which contains no information from $\mathbf{x}_{0}$. The reverse process aims to recover $\mathbf{x}_{0}$ from $\mathbf{x}_{T}$ by learning a parameterized model $p_{\theta}$:
$p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}\big(\bm{\mu}_{\theta}(\mathbf{x}_{t},t),\sigma_{t}^{2}\mathbf{I}\big),$ (2)
where $\bm{\mu}_{\theta}$ is predicted from the noisy sample $\mathbf{x}_{t}$ and the step index $t$. Training minimizes the discrepancy between the true noise and the predicted noise. This framework serves as the foundation for various diffusion-based architectures.
@@
Denoising diffusion probabilistic models (DDPMs) are a class of generative models that create data samples through an iterative denoising process [ddpm]. They work by gradually corrupting real data with noise during training, and then learning to reverse this process. The model is typically trained to predict either the original data or the noise added at various levels of degradation. The loss function is denoted as
[…]
@@
Denoising diffusion implicit models (DDIMs) are a variant of DDPMs that focus on improving generation efficiency [ddim]. While DDPMs rely on stochastic sampling, DDIMs use a deterministic reverse process to generate data, denoted as
@@
Diffusion transformers (DiTs) are used by transforming input images into a sequence of patches, which are then processed by transformer blocks [peebles2023scalable]. To use DiTs, the images need to undergo a patchifying step, which breaks them into smaller tokens. These tokens are then passed through multiple DiT blocks, which use self-attention mechanisms and optional conditioning techniques, such as adaptive layer normalization (adaLN) to handle various input factors like noise timestamp $t$ or class labels. The core computation within a DiT block can be represented as:
$\mathbf{z}^{\prime}=\mathbf{z}+\text{SelfAttention}(\text{adaLN}(\mathbf{z},t,c)),$ (5)
$\mathbf{z}_{\text{out}}=\mathbf{z}^{\prime}+\text{MLP}(\text{adaLN}(\mathbf{z}^{\prime},t,c)),$ (6)
where $\mathbf{z}$ is the input sequence of tokens, and adaLN modulates the network based on the timestep $t$ and an optional class label $c$. This allows the model to effectively learn the denoising function across different noise levels. DiTs can be trained with latent diffusion models to reduce computational cost, and they scale efficiently with increased model size, improving image quality as the model size or computational resources grow. DiT models like Hunyuan-DiT [li2024hunyuandit] have gained attention for their outstanding capability in processing pictures.
2.1.6 Mixture of Experts
Mixture of experts (MoE) is a scalable model architecture that dynamically routes inputs to different specialized subnetworks, known as experts [chen2022towards]. A gating mechanism selects a subset of these experts based on the input, allowing only the chosen experts to process the data. This conditional computation enables models with massive parameter counts to operate efficiently, as only a small portion of the network is active at a time. MoE architectures have demonstrated impressive scalability, particularly in large language models (LLMs) like DeepSeek-V3 [liu2024deepseek], where they improve performance while maintaining inference cost. However, training such models with numerous experts can be challenging, particularly in ensuring the gating mechanism effectively learns to route inputs. In dynamic wireless environments, these challenges manifest as channel conditions and data distributions fluctuate. Recent advancements like DeepEP [deepep2025], DeepGEMM [deepgemm2025] addressed these issue by optimizing MoE training with efficient graphic processing unit (GPU) communication and accelerated matrix operations. In wireless applications, MoEs could enable adaptive processing by activating experts tailored to specific environments, such as urban or rural channel conditions, thereby enhancing model specialization and generalization [wang2025dynamicalmultimodalfusionmixtureofexperts].
@@ 2.2.1 Data Sources and Diversity
Pretraining large models requires diverse datasets. In wireless, this includes channel measurements, CSI matrices, beam logs, and traffic traces collected from real-world networks or high-fidelity simulators. Ensuring coverage across different frequency bands, mobility profiles, and deployment scenarios improves generalization. Unlike conventional pre-training that relies heavily on existing static datasets sourced from the internet, wireless pre-training leverages real-time, multi-modal data collected from a diverse array of internet of things (IoT) devices, including smartphones, sensors, and autonomous as well. This dynamic and heterogeneous data reflects a more accurate and timely representation of the physical world, enhancing the ability of models to understand and predict complex patterns and interactions in various environments. Synthetic data which is generated using ray-tracing, digital twins, or data augmentation strategies to simulate rare or extreme events have also gain consideration [deepmimo]. In semantic communication systems, data must also cover various modalities (text, image, video, audio) and tasks (translation, reconstruction, classification) to ensure the model can learn semantic compression and transmission strategies that generalize across use cases.
@@ 2.2.3 Ethical Considerations
Ethical challenges in constructing large-scale wireless datasets primarily concern two areas: mitigating data bias to ensure fairness, and protecting user privacy from the inherent sensitivity of communication data. Training datasets often reflect existing social, geographic, or behavioral biases; for example, a WLAM trained predominantly on data from urban deployments may underperform in rural conditions, leading to uneven service quality and digital exclusion. Mitigating this risk requires ensuring balanced representation across diverse demographics, devices, and operating environments. In addition to fairness, data privacy is a paramount concern. Technical data such as packet logs, and location traces can inadvertently contain personal identifiers or reveal behavioral patterns, and the broadcast nature of radio signals makes this information particularly vulnerable to misuse. Therefore, all data collection practices must strictly align with regulatory standards like the general data protection regulation [gdpr2016] and the California consumer privacy act [ccpa2018].
@@ 2.3.1 Training Objectives
The choice of training objective plays a critical role in shaping what a model learns during pre-training, with different objectives aligning with specific model architectures and downstream use cases. The next-token prediction objective, commonly used in GPT-style decoder-only models, trains the model to predict the next element in a sequence given the preceding tokens [zhao2023survey]. This approach excels in generative tasks and is well-suited to sequence modeling problems in wireless systems, such as predicting channel evolution or traffic patterns. Masked modeling, used in encoder-only architectures like BERT, involves masking random portions of the input and training the model to reconstruct them using bidirectional context. In wireless applications, this method can be employed for tasks such as recovering missing signal samples or corrupted subcarriers. Sequence-to-sequence generation objectives, typical of encoder-decoder models like T5 or BART, transform an input sequence into a target output, which is particularly useful for end-to-end communication systems, semantic compression, or signal-to-action mappings in adaptive wireless control.
@@ 2.3.2 Distributed and Large-Scale Training
Large model training involves billions of parameters and requires substantial computational resources. Supervised pre-training is typically carried out over distributed systems that involve multiple GPUs connected through high-speed interconnects. To efficiently scale across devices, training workflows employ parallelization strategies such as data parallelism, where each device trains on a separate batch of data with synchronized gradients, or model parallelism, where the model itself is partitioned across devices to handle larger parameter sizes than a single GPU can accommodate.
In pipeline parallelism, the layers of the model are distributed across devices and trained in a staged fashion, enabling higher throughput. WLAM, particularly those involving long input sequences or multimodal inputs, benefit from such techniques. Mixed-precision training reduces memory usage and increases compute efficiency by using lower-precision arithmetic without sacrificing accuracy, and sharded state parallel [fsdp] reduces the overhead of large optimizer memory footprints. Specially designed methods like Dualpipe [dualpipe] also provide an innovative bidirectional pipeline parallelism algorithm and implementation.
@@ 2.4 Reinforcement Learning
Reinforcement learning (RL) provides a powerful framework for WLAM to learn effective policies for decision-making in dynamic wireless environments. The introduction of this framework begins with its fundamental preliminaries, upon which the critical design of reward functions for wireless systems is discussed. The analysis culminates with several advanced algorithms specifically tailored for large AI model training and alignment.
2.4.1 Reinforcement Learning Preliminaries
RL is a computational approach where an agent, such as a network controller or autonomous base station, learns to make optimal decisions by interacting with a dynamic wireless environment. This interaction is formally modeled as a markov decision process, defined by a set of states ($\mathcal{S}$ representing the network status (e.g., channel conditions, traffic load), a set of actions ($\mathcal{A}$) available to the agent (e.g., resource allocation, power control), a reward function ($R$) tied to key performance indicators (e.g., throughput, latency), and state transition dynamics $P(s^{\prime}|s,a)$ that capture the environmental evolution. The decision-making logic of an agent is governed by a policy, $\pi(a|s)$, which specifies the probability of taking action $a$ in state $s$. The primary objective is to learn an optimal policy, $\pi^{*}$, that maximizes the long-term cumulative discounted reward, effectively optimizing network performance over time. This goal is formally expressed as:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}{\pi}\left[\sum{k=0}^{\infty}\gamma^{k}R_{t+k+1}\mid S_{t}=s\right],$ (7)
where $\gamma\in[0,1]$ is the discount factor that balances immediate and future performance. This fundamental objective provides a common mathematical basis for diverse RL algorithms, including the policy-based, value-based, and actor-critic methods discussed later. To achieve (7), RL algorithms often learn a value function to estimate the long-term utility of network states. Unlike supervised learning, which requires static labeled data, RL enables goal-directed optimization in complex, time-varying wireless systems through exploration and exploitation.
@@ 2.4.1 Reinforcement Learning Preliminaries
In the context of large AI models, this RL framework is particularly valuable for tasks where direct supervision is hard to define. It shares strong parallels with adaptive wireless systems, where decisions like resource scheduling must be made under uncertainty with delayed feedback. The dynamic, exploratory nature of RL makes it well-suited for non-stationary or ambiguous objectives, which are common in both advanced wireless networks and large-scale model alignment. By integrating RL with supervised pre-training, models can be fine-tuned to continuously adapt to the evolving demands of dynamic environments. As large models scale, hybrid training strategies could further enhance efficiency. Moreover, by integrating RL with the existing supervised pre-training, models can adjust to the evolving demands of wireless environments.
2.4.2 Reward Function Design
Effective reward function design is paramount for applying reinforcement learning to wireless systems, as it translates high-level performance objectives into a guiding feedback signal for the agent. Adopting a problem-oriented approach, the following analysis categorizes reward design based on the fundamental wireless challenges.
Resource Management and Optimization
The core problem in this domain is the allocation of limited network assets (e.g., bandwidth, power) among competing entities to optimize a system-level utility. The common principle for reward design is to construct a multi-objective utility function, often as a weighted sum of conflicting KPIs, to balance system efficiency with user satisfaction. For example, in dynamic resource allocation for network slicing, the goal is to allocate bandwidth efficiently across different slices. The reward function is designed as a weighted sum of system-wide spectral efficiency (SE) and quality of experience (QoE) [8540003]. While SE is measured at the physical layer, it serves as a crucial system-level feedback for the network-level allocation decision. The objective is to maximize:
$\mathbb{E}\{R(w,d)\}=\mathbb{E}\{\zeta\cdot SE(w,d)+\beta\cdot QoE(w,d)\},$ (8)
where $w$ is the bandwidth sharing solution, $d$ represents the fluctuating demands, and the weights $\zeta$ and $\beta$ balance the importance of spectrum efficiency and user experience.
Mobility Management and Trajectory Optimization
This category focuses on decisions involving physical movement, such as unmanned aerial vehicle (UAV) trajectory planning. The core principle for reward design is to balance the communication performance gains with the physical costs of movement, such as energy consumption or time. The reward is often formulated as an efficiency metric. For instance, in a system where a UAV optimizes its three dimensional trajectory, the reward function is defined as the cumulative sum of the instantaneous energy efficiency achieved at each step of the trajectory [3D-Trajectory]:
$r(s(n),a(n))=\sum_{k=1}^{K}\sum_{n^{\prime}=1}^{n+1}\frac{t_{n^{\prime}}^{u}r_{k,n^{\prime}}}{e_{n^{\prime}}^{uav}}-p_{0},$ (9)
where the term inside the summations represents the instantaneous energy efficiency for user $k$ at a past time slot $n^{\prime}$. Specifically, $t_{n^{\prime}}^{u}$ is the duration of time slot $n^{\prime}$, $r_{k,n^{\prime}}$ is the data rate for ground terminal $k$, and $e_{n^{\prime}}^{uav}$ is the UAV propulsion energy. The double summation aggregates this efficiency metric over all $K$ ground terminals and all time slots from the beginning to the current one. The term $p_{0}$ is a constant penalty applied if the data rate within a time slot falls below a required average, ensuring the overall data delivery task is met. This structure directly incentivizes the agent to learn a trajectory that maximizes energy efficiency throughout its mission.
Security and Adversarial Problems
This category addresses scenarios involving an intelligent adversary, such as an eavesdropper or jammer. The reward design principle is often based on a differential metric that captures the zero-sum nature of the problem. The goal is not just to improve the performance of the legitimate link, but to simultaneously degrade the performance of the adversary. For instance, in physical layer security with cooperative jamming, this principle is applied by maximizing the secrecy rate [he2022deep]. First, the secrecy capacity $C^{s}$ for a single user is explicitly defined as the differential channel capacity between the legitimate user and the eavesdropper:
$C^{s}(u_{k},\alpha_{k})=[C_{\alpha_{k},k}-C^{e}(\alpha_{k})]^{+},$ (10)
where $u_{k}$ is the k-th legitimate user, $\alpha_{k}$ is the index of the access point serving user $u_{k}$, $C_{\alpha_{k},k}$ is the Shannon capacity of the legitimate link, and $C^{e}(\alpha_{k})$ is the capacity of the link to the most effective eavesdropper. The operation $x]^{+}=\max(x,0)$ ensures the secrecy capacity is non-negative. Therefore, the overall reward function is defined as the sum of these individual secrecy capacities across all $K$ users in the system:
$R(S,a)=\sum_{k=1}^{K}C^{s}(u_{k},\alpha_{k}).$ (11)
The action of the agent, denoted by $a$, represents the power allocation vector for all access points. By maximizing this aggregate reward $R(S,a)$, the agent is incentivized to learn a power allocation strategy that intelligently manipulates the signal environment to increase each $C_{\alpha_{k},k}$ while simultaneously suppressing each $C^{e}(\alpha_{k})$, thereby enhancing system-wide security.
Figure 5:   Advanced reinforcement learning algorithms.
2.4.3 Advanced RL Algorithms for Large Model Training and Alignment
The following paragraphs present several advanced algorithms particularly suited to the scale and complexity of large models, as summarized in Fig. 5. The analysis begins with a foundational policy gradient method for stable training, before progressing to sophisticated paradigms designed explicitly for aligning model behavior with complex objectives.
Proximal Policy Optimization
Proximal policy optimization (PPO) [ppo] is a policy gradient RL algorithm that has gained significant popularity due to its stability and efficiency, making it a common choice for fine-tuning LLMs, especially in the context of RL from human feedback. PPO is designed to address the training instability common in earlier policy gradient methods, which are often susceptible to destructively large policy updates that can lead to a sudden collapse in performance. The core idea is to ensure that the updated policy remains proximal, or close, to the policy used for collecting data. To achieve this, PPO maximizes a surrogate objective function that incorporates a clipping mechanism. This mechanism constrains the probability ratio between the new and old policies, effectively discouraging policy updates that are too large. The clipped surrogate objective is denoted as
@@ 2.4.3 Advanced RL Algorithms for Large Model Training and Alignment
which represents the core PPO clipped surrogate objective function $L^{\text{CLIP}}(\theta)$. The expectation $\hat{\mathbb{E}}_{t}$ is over a batch of transitions. $r_{t}(\theta)=\pi_{\theta}(a_{t}|s_{t})/\pi_{\theta_{\text{old}}}(a_{t}|s_{t})$ is the probability ratio of action $a_{t}$ under the current policy $\pi_{\theta}$ versus the policy $\pi_{\theta_{\text{old}}}$ used for data collection. $\hat{A}_{t}$ is the estimated advantage of taking action $a_{t}$ in state $s_{t}$. The clipping mechanism (using hyperparameter $\epsilon$) limits how much the probability ratio can change, enhancing stability. Specifically, when the advantage $\hat{A}_{t}$ is positive, this caps the objective to prevent an overly aggressive policy increase, while for a negative advantage, it sets a bound to avoid excessive penalization. By taking the minimum of the unclipped and clipped objectives, this formulation removes the overall incentive for the policy to change too drastically from its previous version, thereby ensuring more stable and reliable training updates. PPO is widely used due to its simplicity, efficiency, and robustness when scaling to large models.
Reinforcement Learning From Human Feedback
Reinforcement learning from human feedback (RLHF) aligns LLMs with human preferences [rlhf, rlhfpre]. It is especially effective in tasks where human judgment is clear but difficult to formalize. By incorporating human feedback, RLHF helps reduce biases and unsafe behavior, making it central to aligning models like InstructGPT [instructgpt] and OpenAI o1 with human values. Unlike supervised learning, RLHF incorporates human judgment, refining model behavior in tasks where objectives are hard to define. The RLHF pipeline typically involves three stages: supervised fine-tuning on instruction-following examples, training a reward model with human-labeled preferences, and optimizing the model with RL, often using PPO to ensure stability and efficiency, denoted as
@@
which shows the objective maximized during the RL phase of RLHF. The language model policy $\pi_{\theta}$ is optimized to generate responses $y$ for prompts $x$ (from distribution $\mathcal{D}$) that maximize the score from a learned reward model $R_{\phi}(x,y)$, which reflects human preferences. A KL-divergence penalty, weighted by $\beta$, regularizes the policy $\pi_{\theta}$ to stay close to a reference policy $\pi_{\text{ref}}$ (e.g., the initial supervised fine-tuned model), maintaining model capabilities and stability. On the other hand, RLHF faces challenges, including the need for high-quality human feedback, the complexity of reward model training, and computational intensity. Additionally, there are risks of reward hacking [rewardhack], where models exploit reward signals without genuinely improving outputs. Despite these challenges, RLHF remains a powerful method for scalable LLM alignment and has potential applications in WLAM.
@@
Direct Preference Optimization
Direct preference optimization (DPO) [dpo] is a alternative to RLHF that removes the need for an explicit reward model. DPO directly optimizes the language model using human preference data by framing alignment as a classification task. Given pairs of model outputs labeled by human preference, DPO adjusts model parameters to increase the likelihood of preferred responses, typically using a binary cross-entropy loss, as
@@
Group Relative Policy Optimization
@@
which defines the group-relative advantage estimate central to GRPO. For a prompt $x$, the policy generates a group of $K$ candidate responses $\{y_{j}\}_{j=1}^{K}$. The advantage $\hat{A}_{G}$ of a specific response $y_{i}$ is computed as its reward $R(x,y_{i})$ (obtained from a reward model or human feedback) minus the average reward across all $K$ responses in that group. This relative advantage measure replaces the value function estimate used in traditional policy gradient methods like PPO. This approach simplifies training, reduces memory usage, and eliminates the need for large-scale critic gradient updates. Specifically, in DeepSeek-R1-Zero [dsr1], GRPO was applied without any supervised fine-tuning, yielding a model trained purely via RL. While early iterations struggled with readability, later versions like DeepSeek-R1 combined GRPO with multi-stage pretraining and alignment to achieve state-of-the-art results on reasoning tasks. The scalability of GRPO has also influenced other models such as QwQ-32B [qwq32b], which uses a GRPO-inspired two-stage RL method to achieve strong performance with just 32B parameters.